# Efficient Optimization Accelerator Framework for Multi-state Ising Problems

### v1

@@ Conclusion
Current methods employ the brute-force framework to map and solve multistate Ising problems onto conventional Ising machines by using one-hot state encoding. This work presents an alternative approach of vectorized mapping that maps the state using binary encoding. It not only reduces state exploration from to for a q color problem instance but also removes the constraints resulting from the conventional brute-force Ising approach. As a result, the proposed approach converges close to the optimal solution achieved by heuristics and learning based approaches. Exising work [16, 12, 20] on Ising Machines and its derivative frameworks generally aims to achieve accuracy within a certain error range compared to heuristics. However, we show that our method combined with parallel tempering achieves the solution accuracy even better than the heuristics for some of the problems, while retaining the ability to significant speed up on a custom accelerator. We also present a generalized truth table-based method that can be leveraged to map other multi-state problems. Using probabilistic Ising machine architectures, these truth tables to model the neuron interactions can be directly mapped onto higher order multiplexers. We have implemented this architecture on FPGA to benchmark time-to-solution, energy efficiency, and area efficiency. Overall, the hardware implementation consumes 5W power and achieves approximately 1000 - 10000 speedup compared to heuristics on CPU and vectorized mapping on GPU. The presented hardware and software framework provides a new way to substantially expand the capabilities of the Ising machines to accurately handle a wide range of multistate optimization problems in a performance and energy-efficient manner.
@@ 1. Probabilistic Ising Machines
Probabilistic Ising machines follow the principle of the Boltzmann Machine binary neural network. Therefore, the probability distribution corresponding to a given state where is the normalizing/partition function. The distribution is sampled such that the Ising machine stochastically moves towards higher probability states and lower energy. It is enabled by appropriately formulating the update rule which for Hamiltonian of the form , is given by where .
2. Parallel Tempering
The parallel tempering algorithm utilizes multiple Markov chains (replica chains) running at different temperature schedules representing different probability distributions. It allows a broader exploration of the energy landscape and facilitates better mixing to avoid local energy minima solutions. In this work, we implement 100 parallel chains running at a constant temperature geometrically spaced from temp0 (0.01) and tempn (40). After every 15 sampling steps, adjacent pairs of chains are swapped alternatively staring with odd and even chain indexes. The update rule is given as:
3. FPGA Setup PCIe
In this work, we implement the vectorized mapping on the Xilinx Virte UltraScale+ VCU118 FPGA evaluation kit. The FPGA is interfaced with CPU using peripheral component interconnect express (PCIe) interface. In particular, we use open-source xillybus IP core for the interface with data transfer capabilities of 50MB/s. The data is transferred via a memory mapped interface implemented using block memory and a designed memory controller.
The digital implementation of 256 nodes and 16 color accelerators supports a network of 1024 probabilistic Ising nodes. These nodes consist of an LFSR-based pseudorandom number generator with a fixed seed, a lookup table-based sigmoidal activation, higher order multiplexer-based matrix multiplication, and a threshold that generates the output state of the neuron. These states are stored in the local memory of FPGA and then transfered to CPU via PCIe interface.
This work is supported by MURI project from Department of Defense and Office of Naval Research.
C.G. formulated the vectorized scheme, performed the CPU/GPU benchmarks and designed the vectorized accelerator on FPGA. C.G, and S.S co-wrote the manuscript; S.S supervised the research. All authors contributed to discussions and commented on the manuscript.
The authors declare that they have no competing financial interests.
Correspondence and requests for materials can be addressed to either C.G.
(chiraggarg@berkeley.edu) or S.S. (sayeef@berkeley.edu).
Code and Data will be made available on reasonable request by emailing C.G. or S.S.
a, Quadratic unconstrained binary optimization (QUBO) problems are represented as an Ising network with nodes representing the states and edges depicting interactions between them. b, Probabilistic Ising architecture based on the Boltzmann machine update rule to minimize the problem Hamiltonian and energy function. c, The State of the Ising machine evolves and converges to the optimal energy solution. d, Classification of optimization problems based on the state values and problem constraints.
a, The baseline Ising framework employs one-hot encoding to map node color graph coloring problem to nodes. The encoding constraint is enforced by adding an extra term in the Hamiltonian leading to a complex optimization landscape. Since only states are valid among states for a node, this mapping often explores a lot of invalid solution space making it hard for Ising hardware to reach the optimal energy solution. This mapping preserves QUBO Hamiltonian, therefore, the interactions are modeled via binary multiplexers in hardware. b, Vectorized framework maps the color state of a node using hardware nodes. It allows exploration only in valid solution space making it easier for hardware to reach ground-state energy solutions. The solution exploration is further enhanced by the adoption of parallel tempering. This mapping can be modeled in truth table format which is directly mapped to a higher-order multiplexer in hardware.
a, Energy evolution for problem instance mapped and solved using Ising and Vectorized mapping implementation on GPU. (b) Valid and Invalid energy states recorded during solution exploration employing Ising and Vectorized implementation. Ising framework explores completely over invalid solution space making it hard to reach an optimal solution. (c) Vectorized framework combined with parallel tempering enhances the solution exploration avoiding the states being stuck around local minima solution.
a, Physical implementation nodes used in Ising and Vectorized mapping approach for dataset problem [40]. (b) Time-to-solution (TTS) comparison of vectorized mapping implemented on FPGA with Tabucol heuristic on CPU and it equivalent implementation on GPU.
Problem nodes edges colors GNN[33] PI-SAGE[33] Tabucol Probabilistic Ising Vectorized GPU Vectorized FPGA Vectorized+ Parallel Tempering GPU anna 138 493 11 1 0 0 - 0 0 0 myciel5 47 236 6 0 0 0 0 0 0 0 myciel6 95 755 7 0 0 0 0 0 0 0 queen55 25 160 5 0 0 0 0 0 0 0 queen66 36 290 7 4 0 0 - 1 1 0 queen77 49 476 7 15 0 0 - 6 5 0 queen88 64 728 9 7 1 0 - 4 2 1 queen99 81 1056 10 13 1 0 - 5 4 2 queen812 96 1368 12 7 0 0 - 2 2 0 queen1111 121 1980 11 33 17 15 - 20 18 14 queen1313 169 3328 13 40 26 21 - 31 26 21

### v2

@@ Discussion
Current methods employ the one-hot state encoding framework to map and solve multistate spin Ising problems onto Ising machines. This work presents an alternative approach of vectorized mapping that maps the state using binary encoding. It not only reduces state exploration from to for a q color problem instance but also removes the constraints resulting from the constraint-heavy Ising mapping. As a result, the proposed approach converges close to the optimal solution achieved by heuristics and learning based approaches. Existing works [16, 12, 20] on Ising Machines and its derivative frameworks generally aims to achieve accuracy within a certain error range compared to heuristics. However, we show that our method combined with parallel tempering achieves the solution accuracy even better than the heuristics for some of the problems, while retaining the ability to significant speed up on a custom accelerator. We also present a generalized truth table-based method that can be leveraged to map other multi-state problems. Using probabilistic Ising machine architectures, these truth tables to model the neuron interactions can be directly mapped onto higher order multiplexers. We have implemented this architecture on FPGA to benchmark time-to-solution, energy efficiency, and area efficiency. Overall, the hardware implementation consumes 5W power and achieves approximately 10000 speedup compared to Tabucol heuristics and 100000 compared to vectorized mapping on GPU. The presented hardware and software framework provides a new way to substantially expand the capabilities of the Ising machines to accurately handle a wide range of multistate optimization problems in a performance and energy-efficient manner.
1. Parallel Tempering
The parallel tempering algorithm utilizes multiple Markov chains (replica chains) running at different temperature schedules representing different probability distributions. It allows a broader exploration of the energy landscape and facilitates better mixing to avoid local energy minima solutions. In this work, we implement 100 parallel chains running at a constant temperature geometrically spaced from temp0 (0.01) and tempn (40). After every 15 sampling steps, adjacent pairs of chains are swapped alternatively with odd-leading and even-leading chain indexes. The update rule is given as:
2. FPGA Setup PCIe
In this work, we implement the vectorized mapping on the Xilinx Virtex UltraScale+ VCU118 FPGA evaluation kit. The FPGA is interfaced with the CPU using peripheral component interconnect express (PCIe) interface. In particular, we use an open-source xillybus IP core for the interface with data transfer capabilities of 50MB/s. The data is transferred via a memory-mapped interface implemented using block memory and a designed memory controller.
The digital implementation of 256 nodes and 16 color accelerators supports a network of 1024 probabilistic Ising nodes. These nodes consist of an LFSR-based pseudorandom number generator with a fixed seed, a lookup table-based sigmoidal activation, higher order multiplexer-based matrix multiplication, and a threshold that generates the output state of the neuron. These states are stored in the local memory of FPGA and then transferred to CPU via the PCIe interface.
All processed data generated in this study are provided in the main text. The data supporting the plots in this paper can be found in the GitHub repository [48]. Other findings of this study are available from the corresponding authors upon request.
Code will be made available on reasonable request by emailing C.G. or S.S.
Supplementary Information
Efficient Optimization Accelerator Framework for Multi-state Spin Ising Problems
[…]
1Department of Electrical Engineering and Computer Sciences, University of California, Berkeley, CA 94720, USA.
* Corresponding author. Email: chiraggarg@berkeley.edu; sayeef@berkeley.edu
@@ 1. Probabilistic Ising Machines
Probabilistic Ising machines follow the principle of the Boltzmann Machine binary neural network [49]. Therefore, the probability distribution corresponding to a given state where is the normalizing/partition function, represents the states, T denotes temperature coefficient and is Hamiltonian or energy function to be minimized. Ising Hamiltonian is generally of the form , where and represents the state of spins taking binary values {0, 1}, sets the interaction weight between these spins and is bias value for spin . The probability distribution represented by this Ising Hamiltonian is sampled using Gibbs sampling [50] such that the Ising machine stochastically moves toward higher probability states and lower energy. It is enabled by appropriately formulating the update rule, which for Ising Hamiltonian, is given by where .
This work particularly employs single flip Gibbs sampling to update the spins as illustrated in Algorithm 2. For each spin update, the algorithm first calculates the change in Hamiltonian () due to spin flip and compares with a random number from uniform distribution of [0, 1] to update the spin. This comparison predominantly supports the spin flips that cause Hamiltonian to minimize. However, there is still a non-zero probability that it would accept spin updates that cause the Hamiltonian to increase [49, 51, 10]. It gives rise to minor energy fluctuations in energy exploration data reported in Supplementary Fig. S4.
2. Hyperparameter optimization for graph coloring
In order to make a fair comparison between Ising and vectorized mapping, a grid search-based parameter optimization is performed for the parameters in Eq. 1 and in Eq. 2 in the graph coloring Hamiltonian. If the connectivity weighting factor () for this constraint is set relatively higher, the Ising machine may frequently violate the one-hot encoding condition. This can lead to the exploration of invalid solution spaces or failure to enforce correct color assignments, resulting in incorrect graph colorings. On the other hand, if one-hot constraint factor () is too high, the machine may place too much emphasis on satisfying the one-hot constraint, which can hinder the optimization of the actual coloring objective. As a result, it may become trapped in poor local minima where valid but suboptimal solutions prevail. Supplementary Fig. S1 reports the hyperparameter optimization where we report parameter search for each problems separately by running it for 1000 iteration steps at = 0.2 and reporting the coloring error (incorrectly colored edges divided by total edges). The optimal settings from Supplementary Fig. S1 are chosen in this work for testing graph coloring problem instances using the probabilistic Ising framework. The simulated bifurcation framework[44] used in this work only accepts the Ising Hamiltonian to minimize. Accordingly, the parameters in Eq. 1 and in Eq. 2 are optimized by running the framework for 10000 iteration steps. The optimal values identified in Supplementary Fig. S2 are then used to construct the Hamiltonian, which is subsequently solved using the simulated bifurcation method for the graph coloring problem. For vectorized mapping implementation, only temperature effects the results and is chosen equal to 0.2 for all problem instances.
3. Vectorized Mapping Benchmark for Citation Graph Dataset
To demonstrate the scalability of vectorized mapping approach, we benchmark the citation datasets (Cora [52], Citeseer [53], and Pubmed [54]). These problems are often used for graph-based benchmark experiments and have also been used for testing graph-coloring based algorithms [33, 34]. Table 2 confirms that the solution accuracy advantage of the proposed vectorized mapping holds for large size graph problems. Supplementary Fig. S5 shows the time-to-solution for these citation graphs. It confirms that the vectorized mapping on GPU takes around one order of magnitude more time compared to the Tabucol heuristics aligning with the scaling trends in Fig. 4c. Moreover, in probabilistic Ising hardware, the overall hardware scaling is primarily influenced by the number of physical nodes required to solve a problem [55, 12]. As a result, it follows the scaling behavior of physical nodes with respect to problem size illustrated in Fig. 4a.
4. Traveling Salesman Problem
The Traveling Salesman Problem (TSP) is a classic NP-hard optimization problem that aims to determine the shortest possible route in which a given set of cities is visited exactly once. Owing to its computational complexity and broad applicability in areas such as logistics [56], circuit design[2], and DNA sequencing[57], TSP remains a central focus in combinatorial optimization research. Recently, Ising machines are being explored to tackle this problem [31, 58]. To solve -cities problems, these solvers require spins modeled via a lattice-like graph representing city number and spin. The Hamiltonian is given as follows:
where is the connectivity weight factor, one-hot constraint factor, represents weight defined as distance between city k and l, and is the spin value for visiting city at position in the tour. The first term in the Hamiltonian represents the total weight for given values of spins. However, the second and third terms are to enforce one-hot encoding constraints that penalize assigning multiple cities for visiting at the same position and visiting a city multiple times in a tour, respectively. In this work, we use the problem instance from the tsplib dataset [59] and use the first subset of the weight matrix to generate smaller problem instances. The weight matrix is normalized with the maximum distance in it [58]. For the Ising framework, we employ Gibbs Sampling described in Algorithm 2 and choose temperature value 0.02. Further, we exploit the grid-search hyperparameter optimization for parameters and as shown in Supplementary Fig. S6. Each problem instance is run for 4000 iteration steps, and optimized for the optimality gap. It is given by the Supplementary Eq. S2, where the optimal solution is calculated using the Lin-Kernighan (LK) heuristics [60].
We propose the vectorized mapping framework to tackle the TSP problem, which maps city problem to . A spin vector represents the position of the city in the tour and is defined as where . Supplementary Algorithm A1 describes the formulation of vectorized mapping Hamiltonian and operator function . If and are consecutive cities, the Hamiltonian is incentivized to optimize for the tour cost. However, it is penalized when two cities are visited at the same position in the tour guided by the hyperparameter . Further, the formulation also restricts to be within its range of N. For the vectorized mapping framework, we employ the Gibbs sampling described in Algorithm 2 and choose the temperature value 0.2. The hyperparameter is optimized for the optimality gap, as shown in Supplementary Fig. S7.
To evaluate the broader applicability of the proposed vectorized mapping framework, we conducted additional experiments on the Traveling Salesman Problem (TSP). Both the Ising model and the proposed vectorized mapping were implemented on Nvidia A100 Tensor Core GPU and tested across TSP instances involving up to 14 cities. Each instance was solved 500 times, involving 4000 iteration or update steps per run. As shown in Supplementary Fig. S8, the proposed approach achieves higher success probabilities and shorter time-to-solution compared to the Ising model. Additionally, it consistently produces tours that are closer to those found by the well-established Lin–Kernighan (LK) heuristic, indicating improved convergence toward high-quality solutions (see Supplementary Fig. S9). These results suggest that the vectorized mapping framework is an effective strategy for addressing complex combinatorial optimization problems beyond graph coloring, such as TSP.
[60].